At the beginning of a basketball season, the Panthers won 33 games and lost 55 games. At this rate, how many games will they win in a normal 100-game season?
Answer by stanbon(75887)

At the beginning of a basketball season, the Panthers won 33 games and lost 55 games. At this rate, how many games will they win in a normal 100-game season?
-----------------
Use a proportion:
wins/total = wins/total
-------
x/100 = 33/(33+55)
-----
x/100 = 33/88
x = 100(3/8) = 37.5 games (# they will win in 100-game season)
---------------------
Cheers,
Stan H.
=================